determine inverse function of the given functions y=5x+4, y=ln(3x)

Answer by tommyt3rd(5050)

y=5x+4
x=(y-4)/5 so inverse is...
y=1/5 x - 4/5



y=ln(3x)
x=1/3 e^y so inverse is...
y=1/3 e^x